An investment will pay you $28,500 in 12 years. The stated interest rate is 7% (APR). If interest is compounded semi-annually, what is the present value? Enter your response below rounded to two decimal places.
Answer #1

Answer: $ 12,482

n = 12 x 2 = 24

i = 7 % x 1/2 = 3.5 %

PV 3.5%, n = 24 = ( 1 / 1.035 ) 24 = 0.43796

Present value of the investment = $ 28,500 * 0.43796 = $ 12,481.86